OpenStudy (anonymous):

Use the quadratic formula to solve the equation. 2x^2 -3x= -6 The solution set is { }

OpenStudy (anonymous):

\[2x^2 - 3x + 6 =0 \] \[x = \frac{ 3 \pm \sqrt{9 - 4*6*2}}{4}\] \[x = \frac{3 \pm i\sqrt{39}}{4}\] Solution is not real.